# Computing the Matching Distance of $2$-Parameter Persistence Modules from Critical Values

### v2

@@ 3.1 Preliminary claims
Lemma 3.2.
Proof.
More explicitly, the multi-bijection $\gamma_{M}$ behaves as follows.
Similarly, there is some $j\in\{1,2\}$ such that
For the rest of the section, we fix a line $L$ with positive slope, we let $P\subseteq\operatorname{dgm}M^{L}$ and $Q\subseteq\operatorname{dgm}N^{L}$, and we take $\sigma\colon P\rightarrow Q$ to be a matching such that the pair $u,v\in\overline{C}_{M}\cup\overline{C}_{N}$ achieves the cost of $\sigma$: ${\rm cost}(\sigma)=\hat{m}^{L}\frac{|p_{L}(u)-p_{L}(v)|}{\delta},$ with
$\delta=\begin{cases}1,&\text{if ${\rm cost}(\sigma)$ is achieved by a matched pair, }\\ 2,&\text{if ${\rm cost}(\sigma)$ is achieved by an unmatched point .}\end{cases}$
@@
      Suppose that $\{1\}\subseteq A^{L}_{\mathrm{l.u.b.}(u,v)}$ and $v_{2}\geq r_{2}$. By Equation 7, taking $m_{1}^{L^{\prime}}<m_{1}$ rotates $L$ away from the diagonal direction and does not decrease the cost of the matching $\sigma$. In this case, we may hit no point of $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ before we get to the vertical line through $r$, that is, we may eventually hit the point $[0:0:1]\in\ell_{\infty}$.
@@
      Now, suppose that $\{1\}\subseteq A^{L}_{\mathrm{l.u.b.}(u,v)}$ and $v_{2}<r_{2}$. This time, we take $m_{1}^{L^{\prime}}>m_{1}$, which rotates the line $L$ towards the diagonal direction without decreasing its cost. In this case, we may hit no point of $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ before we get to the diagonal line through $r$, that is, we may eventually hit the point $[0:1:1]\in\overline{\Omega}$.
@@
When the slope of $L$ is smaller than $1$, that is $\hat{m}^{L}=m_{2}<m_{1}=1$, the argument is very similar except that for the case 2(c) where we may hit the point $[0:1:0]\in\ell_{\infty}$ instead of the point $[0:0:1]$. The case when the slope of $L$ is exactly 1 is excluded by the assumption that $L$ intersects $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ only at the proper point $r$. ∎
@@ 3.2 Vertical and Horizontal Lines
$\begin{split}\varepsilon_{\infty}:=&\min\{m_{1}\in\mathbb{R}\mid[0:m_{1}:m_{2}]\in\overline{\Omega},m_{2}=1\}\ ,\\ c_{min}:=&{\mathrm{g.l.b.}}\left(\overline{C}_{M}\cup\overline{C}_{N}\cup\Omega_{P}\right)\ ,\\ c_{max}:=&{\mathrm{l.u.b.}}\left(\overline{C}_{M}\cup\overline{C}_{N}\cup\Omega_{P}\right)\ ,\\ \varepsilon:=&\min\{|x_{1}-y_{1}|\mid x_{1}\neq y_{1}\,,\,x,y\in\overline{C}_{M}\cup\overline{C}_{N}\cup\Omega_{P}\}\ .\\ \end{split}$ (9)
Assume, without loss of generality, that $c_{2}<c^{\prime}_{2}$. Taking $\varepsilon/0$ to be equal to +$\infty$+ ∞, choose
@@
Let $V\subseteq\mathbb{R}^{2}$ be a vertical line, and let $c,c^{\prime}$ be distinct points on $V$ such that the line segment between them does not contain any point of $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$. Let $L$ be a line through a point $r\in V$ between $c$ and $c^{\prime}$ with direction $\vec{m}=(m_{1},1)$, $10<m_{1}<$ 1. Then the limit of the weighted bottleneck distance along $L$ as it rotates around $r$ toward the vertical direction, written as
@@
We consider the behavior of the weighted bottleneck distance along lines $L(r,m_{1})$ with direction $\vec{m}=(m_{1},1)$, and $10<m_{1}<$ 1, through a point $r$ on the open segment of $V$ between $c$ and $c^{\prime}$, as $m_{1}$ tends to 00.
First of all we observe that, for any two points $r$ and $r^{\prime}$ on such open segment, and for sufficiently small $m_{1}>0$, Lemma 3.10 ensures that $L:=L(r,m_{1})\sim_{\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}}L(r^{\prime},m_{1})=:L^{\prime}$. Hence, by Property (Q’), if the optimal matching achieving the bottleneck distance along $L$ is $\sigma$, then the optimal matching achieving the bottleneck distance along $L^{\prime}$ is $\Gamma_{L,L^{\prime}}(\sigma)$, and that the matched pair giving the cost of $\Gamma_{L,L^{\prime}}(\sigma)$ is the image of the matched pair giving the cost of $\sigma$ under $\Gamma_{L,L^{\prime}}$: If $p_{L}(u),p_{L}(v)$ give the bottleneck distance along $L$, then $p_{L^{\prime}}(u),p_{L^{\prime}}(v)$ give the bottleneck distance along $L^{\prime}$. In other words, for any line $L(\cdot,m_{1})$ that is steep enough and intersects $V$ between $c$ and $c^{\prime}$, the bottleneck distance is determined by the same pair $u,v$ of critical points.
@@
    If both $A_{v}^{L(r,m_{1})}$ and $A_{u}^{L(r,m_{1})}$ contain $2={\operatorname*{\arg\!\max}}_{i\in\{1,2\}}\{m_{i}\}$, then the weighted bottleneck distance is $m_{1}\frac{|u_{2}-v_{2}|}{\delta}$, which is independent of $r$ and tends to 00 when $m_{1}$ tends to 00. In particular, for any $r\in V$ between $c$ and $c^{\prime}$ and any $0<m_{1}<\frac{\epsilon}{|u_{2}-v_{2}|}$, we have that $m_{1}\frac{|u_{2}-v_{2}|}{\delta}<\epsilon$, proving the uniform convergence of $\hat{m}^{L(r,m_{1})}d_{B}\left(\operatorname{dgm}M^{L(r,m_{1})},\operatorname{dgm}N^{L(r,m_{1})}\right)$ to 00 as $m_{1}$ tends to 00 in this case.
@@
    If $A_{v}^{L(r,m_{1})}$ and $A_{u}^{L(r,m_{1})}$ contain $1={\operatorname*{\arg\!\min}}_{i\in\{1,2\}}\{m_{i}\}$, then the weighted bottleneck distance is constantly equal to $\frac{|u_{1}-v_{1}|}{\delta}$, independently of both $r$ and $m_{1}$. Hence, it converges uniformly (for steep enough lines intersecting $V$ between $c$ and $c^{\prime}$) to $\frac{|u_{1}-v_{1}|}{\delta}$ as $m_{1}$ tends to 00 in this case.
@@
    If $A_{u}^{L(r,m_{1})}=\{1\}$ and $A_{v}^{L(r,m_{1})}=\{2\}$, and $p_{L(r,m_{1})}(u)-p_{L(r,m_{1})}(v)\leq 0$, then the weighted bottleneck distance is $\frac{1}{\delta}\cdot\left(m_{1}(v_{2}-r_{2})+r_{1}-u_{1}\right)$. As $m_{1}$ tends to 00, this tends to $\frac{1}{\delta}\cdot(r_{1}-u_{1})$ regardless of $r$ (that is, regardless of $r_{2}$, as $r_{1}$ is the same for any point $r$ between $c$ and $c^{\prime}$), thanks to the fact that $r_{2}$ is bounded between $c_{2}$ and $c^{\prime}_{2}$.
@@
    Therefore, for any $r\in V$ between $c$ and $c^{\prime}$ and any $0<m_{1}<\frac{\epsilon}{\max\{|v_{2}-c_{2}|,|v_{2}-c_{2}^{\prime}|\}}$, we have $\left|m_{1}\cdot\frac{|u_{2}-v_{2}|}{\delta}-\frac{1}{\delta}\cdot(r_{1}-u_{1})\right|<\epsilon$, proving uniform convergence of the weighted bottleneck distance along $L(r,m_{1})$ to $\frac{1}{\delta}\cdot(r_{1}-u_{1})$ as $m_{1}$ tends to 00 in this case.
@@
    If $A_{u}^{L(r,m_{1})}=\{1\}$ and $A_{v}^{L(r,m_{1})}=\{2\}$, and $p_{L(r,m_{1})}(u)-p_{L(r,m_{1})}(v)\geq 0$, then, by an analogous argument to the one made in the previous case, we can see that the weighted bottleneck distance along $L(r,m_{1})$ converges uniformly to $\frac{1}{\delta}\cdot(u_{1}-r_{1})$ as $m_{1}$ tends to 00 in this case.
@@
of the weighted bottleneck distance of the line $L$ through the point $r\in V$ with direction $\vec{m}=(m_{1},1)$, $10<m_{1}<$ 1, as $L$ rotates toward the vertical direction, does not depend on the point $r$.
@@
Let us now assume that $\hat{m}^{L(r)}d_{B}(\operatorname{dgm}M^{L(r)},\operatorname{dgm}N^{L(r)})$ tends to the value $\ell$, which has to be the same for all $r$ sufficiently close to $c$ by Lemma 3.11, and $\hat{m}^{L(c)}d_{B}(\operatorname{dgm}M^{L(c)},\operatorname{dgm}N^{L(c)})$ tends to the value $\ell(c)$, as $L(c)$ and $L(r)$ tend to $V$ by rotating counterclockwise about $r$ and $c$, respectively. Let us also assume, by contradiction, that $|\ell(c)-\ell|>0$ and take $30<\varepsilon<|\ell(c)-\ell|/$ 3.
By internal stability [23, Thm. 2], we know that the weighted bottleneck distance for $L(r)$ tends to that of $L(c)$ as $r$ tends to $c$, so
@@ 3.3 Proof of the main theorem
We are now ready to prove Theorem 3.1, thanks to Proposition 2.10, Lemma 3.4, 3.7, and Corollary 3.14, which we restate below for convenience.
@@
If $\operatorname{dgm}(M^{L}),\operatorname{dgm}(N^{L})$ have a different number of points at infinity, then $d_{B}\left(\operatorname{dgm}(M^{L}),\operatorname{dgm}(N^{L})\right)=\infty$. For any line $L^{\prime}\sim_{\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}}L$, the points at infinity of $\operatorname{dgm}(M^{L})$ (resp. $\operatorname{dgm}(N^{L})$) are in bijection with those of $\operatorname{dgm}(M^{L^{\prime}})$ (resp. $\operatorname{dgm}(N^{L^{\prime}})$) via $\gamma_{M}$ (resp. $\gamma_{N}$) as defined in Lemma 3.2. So, $\operatorname{dgm}(M^{L^{\prime}})$ and $\operatorname{dgm}(N^{L^{\prime}})$ will also have a different number of points at infinity, implying that $d_{B}\left(\operatorname{dgm}(M^{L^{\prime}}),\operatorname{dgm}(N^{L^{\prime}}\right)=\infty$ as well. By internal stability [23, Thm. 2], the bottleneck distance is equal to infinity on any line on the closure [^7] Here, two lines $L$ and $L^{\prime}$ endowed with the standard parametrization $u=\vec{m}\cdot t+b$ and $u=\vec{m}^{\prime}\cdot t+b^{\prime}$, respectively, are $\varepsilon$-distant if $\max\{\|\vec{m}-\vec{m}^{\prime}\|_{\infty},\|b-b^{\prime}\|_{\infty}\}=\varepsilon$. of the equivalence class of $L$. In this closure, we can always find a line passing through two points, one in $\overline{C}_{M}\cup\overline{C}_{N}\cup\Omega_{P}$ and the other one in $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$, implying the claim.

### v3

@@ 3.1 Preliminary claims
In this section, we fix a line $L$ with positive slope, we let $P\subseteq\operatorname{dgm}M^{L}$ and $Q\subseteq\operatorname{dgm}N^{L}$, and we take $\sigma\colon P\rightarrow Q$ to be a matching such that the pair $u,v\in\overline{C}_{M}\cup\overline{C}_{N}$ achieves the cost of $\sigma$: ${\rm cost}(\sigma)=\hat{m}^{L}\frac{|p_{L}(u)-p_{L}(v)|}{\delta},$ with
[…]
@@
      Suppose that $\{1\}\subseteq A^{L}_{\mathrm{l.u.b.}(u,v)}$ and $v_{2}\geq r_{2}$. By Equation 7, taking $m_{1}^{L^{\prime}}<m_{1}$ rotates $L$ away from the diagonal direction and does not decrease the cost of the matching $\sigma$. In this case, we may hit no point of $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ before we get to the vertical line through $r$, that is, we may eventually hit the point $0:0:1]\in\ell_{\infty}$.
@@
      Now, suppose that $\{1\}\subseteq A^{L}_{\mathrm{l.u.b.}(u,v)}$ and $v_{2}<r_{2}$. This time, we take $m_{1}^{L^{\prime}}>m_{1}$, which rotates the line $L$ towards the diagonal direction without decreasing its cost. In this case, we may hit no point of $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ before we get to the diagonal line through $r$, that is, we may eventually hit the point $0:1:1]\in\overline{\Omega}$.
@@
When the slope of $L$ is smaller than $1$, that is $\hat{m}^{L}=m_{2}<m_{1}=1$, the argument is very similar except that for the case 2(c) where we may hit the point $0:1:0]\in\ell_{\infty}$ instead of the point $[0:0:1]$. The case when the slope of $L$ is exactly 1 is excluded by the assumption that $L$ intersects $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ only at the proper point $r$. ∎
@@ 3.2 Vertical and Horizontal Lines
[…]
Assume, without loss of generality, that $c_{2}<c^{\prime}_{2}$. Taking $\varepsilon/0$ to be equal to $\infty$, choose
@@
Let $V\subseteq\mathbb{R}^{2}$ be a vertical line, and let $c,c^{\prime}$ be distinct points on $V$ such that the line segment between them does not contain any point of $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$. Let $L$ be a line through a point $r\in V$ between $c$ and $c^{\prime}$ with direction $\vec{m}=(m_{1},1)$, $0<m_{1}<1$. Then the limit of the weighted bottleneck distance along $L$ as it rotates around $r$ toward the vertical direction, written as
@@
We consider the behavior of the weighted bottleneck distance along lines $L(r,m_{1})$ with direction $\vec{m}=(m_{1},1)$, and $0<m_{1}<1$, through a point $r$ on the open segment of $V$ between $c$ and $c^{\prime}$, as $m_{1}$ tends to 0.
First of all we observe that, for any two points $r$ and $r^{\prime}$ on such open segment, and for sufficiently small $m_{1}>0$, Lemma 3.9 ensures that $L:=L(r,m_{1})\sim_{\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}}L(r^{\prime},m_{1})=:L^{\prime}$. Hence, by Property (Q’), if the optimal matching achieving the bottleneck distance along $L$ is $\sigma$, then the optimal matching achieving the bottleneck distance along $L^{\prime}$ is $\Gamma_{L,L^{\prime}}(\sigma)$, and that the matched pair giving the cost of $\Gamma_{L,L^{\prime}}(\sigma)$ is the image of the matched pair giving the cost of $\sigma$ under $\Gamma_{L,L^{\prime}}$: If $p_{L}(u),p_{L}(v)$ give the bottleneck distance along $L$, then $p_{L^{\prime}}(u),p_{L^{\prime}}(v)$ give the bottleneck distance along $L^{\prime}$. In other words, for any line $L(\cdot,m_{1})$ that is steep enough and intersects $V$ between $c$ and $c^{\prime}$, the bottleneck distance is determined by the same pair $u,v$ of critical values.
@@
    If both $A_{v}^{L(r,m_{1})}$ and $A_{u}^{L(r,m_{1})}$ contain $2={\operatorname*{\arg\!\max}}_{i\in\{1,2\}}\{m_{i}\}$, then the weighted bottleneck distance is $m_{1}\frac{|u_{2}-v_{2}|}{\delta}$, which is independent of $r$ and tends to 0 when $m_{1}$ tends to 0. In particular, for any $r\in V$ between $c$ and $c^{\prime}$ and any $0<m_{1}<\frac{\epsilon}{|u_{2}-v_{2}|}$, we have that $m_{1}\frac{|u_{2}-v_{2}|}{\delta}<\epsilon$, proving the uniform convergence of $\hat{m}^{L(r,m_{1})}d_{B}\left(\operatorname{dgm}M^{L(r,m_{1})},\operatorname{dgm}N^{L(r,m_{1})}\right)$ to 0 as $m_{1}$ tends to 0 in this case.
@@
    If $A_{v}^{L(r,m_{1})}$ and $A_{u}^{L(r,m_{1})}$ contain $1={\operatorname*{\arg\!\min}}_{i\in\{1,2\}}\{m_{i}\}$, then the weighted bottleneck distance is constantly equal to $\frac{|u_{1}-v_{1}|}{\delta}$, independently of both $r$ and $m_{1}$. Hence, it converges uniformly (for steep enough lines intersecting $V$ between $c$ and $c^{\prime}$) to $\frac{|u_{1}-v_{1}|}{\delta}$ as $m_{1}$ tends to 0 in this case.
@@
    If $A_{u}^{L(r,m_{1})}=\{1\}$ and $A_{v}^{L(r,m_{1})}=\{2\}$, and $p_{L(r,m_{1})}(u)-p_{L(r,m_{1})}(v)\leq 0$, then the weighted bottleneck distance is $\frac{1}{\delta}\cdot\left(m_{1}(v_{2}-r_{2})+r_{1}-u_{1}\right)$. As $m_{1}$ tends to 0, this tends to $\frac{1}{\delta}\cdot(r_{1}-u_{1})$ regardless of $r$ (that is, regardless of $r_{2}$, as $r_{1}$ is the same for any point $r$ between $c$ and $c^{\prime}$), thanks to the fact that $r_{2}$ is bounded between $c_{2}$ and $c^{\prime}_{2}$.
@@
    Therefore, for any $r\in V$ between $c$ and $c^{\prime}$ and any $0<m_{1}<\frac{\epsilon}{\max\{|v_{2}-c_{2}|,|v_{2}-c_{2}^{\prime}|\}}$, we have $\left|m_{1}\cdot\frac{|u_{2}-v_{2}|}{\delta}-\frac{1}{\delta}\cdot(r_{1}-u_{1})\right|<\epsilon$, proving uniform convergence of the weighted bottleneck distance along $L(r,m_{1})$ to $\frac{1}{\delta}\cdot(r_{1}-u_{1})$ as $m_{1}$ tends to 0 in this case.
@@
    If $A_{u}^{L(r,m_{1})}=\{1\}$ and $A_{v}^{L(r,m_{1})}=\{2\}$, and $p_{L(r,m_{1})}(u)-p_{L(r,m_{1})}(v)\geq 0$, then, by an analogous argument to the one made in the previous case, we can see that the weighted bottleneck distance along $L(r,m_{1})$ converges uniformly to $\frac{1}{\delta}\cdot(u_{1}-r_{1})$ as $m_{1}$ tends to 0 in this case.
@@
of the weighted bottleneck distance of the line $L$ through the point $r\in V$ with direction $\vec{m}=(m_{1},1)$, $0<m_{1}<1$, as $L$ rotates toward the vertical direction, does not depend on the point $r$.
@@
Let us now assume that $\hat{m}^{L(r)}d_{B}(\operatorname{dgm}M^{L(r)},\operatorname{dgm}N^{L(r)})$ tends to the value $\ell$, which has to be the same for all $r$ sufficiently close to $c$ by Lemma 3.10, and $\hat{m}^{L(c)}d_{B}(\operatorname{dgm}M^{L(c)},\operatorname{dgm}N^{L(c)})$ tends to the value $\ell(c)$, as $L(c)$ and $L(r)$ tend to $V$ by rotating counterclockwise about $r$ and $c$, respectively. Let us also assume, by contradiction, that $|\ell(c)-\ell|>0$ and take $0<\varepsilon<|\ell(c)-\ell|/3$.
By internal stability [28, Thm. 2], we know that the weighted bottleneck distance for $L(r)$ tends to that of $L(c)$ as $r$ tends to $c$, so
@@ 3.3 Proof of the main theorem
We are now ready to prove Theorem 3.1, thanks to Proposition 2.11, Lemma 3.3, 3.6, and Corollary 3.13, which we restate below for convenience.
@@
If $\operatorname{dgm}(M^{L}),\operatorname{dgm}(N^{L})$ have a different number of points at infinity, then $d_{B}\left(\operatorname{dgm}(M^{L}),\operatorname{dgm}(N^{L})\right)=\infty$. For any line $L^{\prime}\sim_{\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}}L$, the points at infinity of $\operatorname{dgm}(M^{L})$ (resp. $\operatorname{dgm}(N^{L})$) are in bijection with those of $\operatorname{dgm}(M^{L^{\prime}})$ (resp. $\operatorname{dgm}(N^{L^{\prime}})$) via $\gamma_{M}$ (resp. $\gamma_{N}$) as defined in Lemma 2.10. So, $\operatorname{dgm}(M^{L^{\prime}})$ and $\operatorname{dgm}(N^{L^{\prime}})$ will also have a different number of points at infinity, implying that $d_{B}\left(\operatorname{dgm}(M^{L^{\prime}}),\operatorname{dgm}(N^{L^{\prime}}\right)=\infty$ as well. By internal stability [28, Thm. 2], the bottleneck distance is equal to infinity on any line on the closure [^7] Here, two lines $L$ and $L^{\prime}$ endowed with the standard parametrization $u=\vec{m}\cdot t+b$ and $u=\vec{m}^{\prime}\cdot t+b^{\prime}$, respectively, are $\varepsilon$-distant if $\max\{\|\vec{m}-\vec{m}^{\prime}\|_{\infty},\|b-b^{\prime}\|_{\infty}\}=\varepsilon$. of the equivalence class of $L$. In this closure, we can always find a line passing through two points, one in $\overline{C}_{M}\cup\overline{C}_{N}\cup\Omega_{P}$ and the other one in $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$, implying the claim.
@@
3.4 Connections to related works
We highlight the relationship of the contents of Section 3.1 -3.3 to both the topological approach papers [15, 18], and the dual approach papers [25, 7].
Within the topological approach, Lemma 3.5 corresponds to Theorem 4 from [15]; both results state when it is possible to continuously move a line through parameter space without decreasing the cost of any matching. For us, this result is integral in proving that the matching distance is achieved as a maximum, rather than a supremum. Differently, in [15], it is used to prove this realization as a maximum for the coherent matching distance, a modification of the matching distance restricted to the closure of an open connected subset of lines not passing through a switch point (rephrased in our language).
The use of vertical and horizontal lines in biparametric persistence, as well as the extension of the bottleneck distance to these lines, is discussed in [18] Section 4, albeit from a different perspective. Notably, Lemma 3.10, Theorem 3.11, and Definition 3.12 are related to the approach described in [18].
Passing to the comparison with [18], which appeared on arXiv at the same time as this manuscript (cf. [2]), we observe that the set of switch points $\Omega$ in Theorem 3.1 is closely related to the set of special values introduced in [18] (Definition 5.1). Lines passing through a switch point $\omega$ correspond to the lines $r(a,b)$ associated with a “special value” $(a,b)$ in the topological case.
Moreover, the statement of Theorem 3.1 corresponds to Theorem 5.4 in [18], although the analogous theorem in the topological case does not exclude vertical and horizontal lines. The proofs of these theorems have similarities as well; a crucial step in proving Theorem 5.4 involves showing that the line $L$ can be continuously moved without reducing the optimal matching cost until it reaches a line of slope $0,1,\infty,$ or a line associated with a special value.
We next highlight similarities and differences of the results of Section 3.3 and [25].
In [25], the line arrangement $\mathcal{A}_{0}$ consists of lines in the dual space which correspond to parameterizations of lines in the primal space through critical values. This line arrangement is further refined to $\mathcal{A}$ (Definition 7), with the goal of obtaining a finite polyhedral decomposition of the dual space for which one may determine a “consistent matching” for computing the weighted bottleneck distance for all lines in a given polyhedron (Theorem 9).
Analogously, and in the primal space, the preliminary claims in Section 2 are all used to show, in the proof of Theorem 3.1, that one may similarly find a “consistent matching” for computing the weighted bottleneck distance for all lines in a given equivalence class, once the equivalence classes are generated by the set of critical values and switch points. In particular, the refinement of the arrangement $\mathcal{A}_{0}$ to $\mathcal{A}$ mirrors the refinement of the equivalence relation achieved when adding the set of switch points to the set of critical values and the diagonal direction (the line $m=1$ in the dual space is also added to $\mathcal{A}$). More about the difference between our switch points and the lines of the arrangement $\mathcal{A}$ will be discussed in Section 4.1.
Following [25], they are able to show that, on any face in the line arrangement, the weighted bottleneck distance obtains a supremum at a boundary vertex of this face, or as the limit of an outer segment (Lemma 10). The latter involves considering horizontal lines (in the primal plane), which are excluded by the approach followed here.
In summary, the main differences between [25, Lemma 10] and our Theorem 3.1, apart from the use of point-line duality, are:
  • Theorem 3.1 states that the matching distance is achieved as a maximum, while in [25, Lemma 10] it is achieved as a supremum (however easy to compute).
  • Theorem 3.1 avoids considering horizontal lines, i.e. points $(m,q)$ with $m=0$ in the dual space (vertical lines are implicitly not needed also in [25]).
Likewise, in [7] the initial line arrangement of [25] is refined but based on a real parameter $\lambda>0$: this time the crucial property is that within each region in the dual plane determined by the line arrangement, the weighted bottleneck distance is either $\leq\lambda$ or $\lambda$.